OpenStudy (anonymous):

Multiply x to the 1 third power times x to the 3 sevenths power.

OpenStudy (nurali):

\[x ^{\frac{ 1 }{ 3 }}*x ^{\frac{ 3 }{ 7 }}\] \[x ^{\frac{ 1 }{ 3 }+\frac{ 3 }{ 7 }}\] \[x ^{\frac{ 16 }{ 21 }}\]